shift summary for evening 7/3/99

Theodoros Alexopoulos (theoalex@fnal.gov)
Sun, 4 Jul 1999 00:02:12 -0500 (CDT)

Shift Summary for evening shift, 7/3/99

On Shift: AS, TA
Other People present: EB via phone

Hours of beam available at all intensities: ~7.5h
Hours of beam available at nominal intensity: ~7.5h
Hours of E832 data (special runs not included): ~6.0h
Hours of special runs (laser scans, muon runs...):~0.2h

Accounting of lost time:
starting/stopping runs, controlled access to turn MU3X power supply off.

========================================================================
Runs ended this shift:
Run 13802 - 16M events - BEAM_832
Run 13803 - 5.9M events - BEAM_832
Run 13804 - 0.3M events - BEAM_MU_4E12

Tapes brought to FCC this shift:
QKC079,QKC080,QKC081,QKC082
QKS079,QKS080,QKS081,QKS082
QKK079,QKK080,QKK081,QKK082
QKB028,QKB029
QKN005,QKN006, QKNL05,QKNL06
QKLA05,QKSM03,QKPE04,QKCC03,QKMU08,QKCR03

========================================================================
Accesses during shift: to turn the MU3X fluke-COW power supply off
(this supplies HV to channels 21-->40)

========================================================================
Overview of shift:
Smooth BEAM_832 datataking, apart from the short controlled access

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13802 16:00 run 13802 BEAM_832_NOKTEV1 in progress
17:05 13802 is over. We switch to BEAM_832 with ktev1.
Loading new tapes, the tape drive tps2d2 on ktev2
gives errors (all its lights are flashing).
We tried to clean it or to load new tape, with no
luck.
Paged Ed, he modified the BEAM_832 tape
configuration file to use tps3d2 tape drive instead.
13803 1 17:45 run starts successfully, with ktev1 working fine.
19:00 sdaq MU3X error indicates that some of the MU3X
channels (21-->40) are as high as -3000V!.
The MU3X fluke-COW power supply can't regulate the HV.
The MU3X illumination online plots reflects this
problem as well.
Trying to get RT on the phone, with no success.
21:00 After a long discussion with Ed, AS and TA go to
controlled access. Recycling the front HV and AC
switches of the power supply,
doesn't fix it. Actually this
power supply can't be turned on anymore. It was
left off.
We took the access during the NM4AN switching
polarity procedure, so didn't really loose any time.
13804 21:30 Do a muon run.
Switch NM4AN to positive.
13805 21:55 Start BEAM_832 run, with half of the MU3X
counters off.
22:35 Intensity drops to 2.5E12 for 5 spills.
Call MCR, they do MI tuning, but they bring
intensity back to 5E12.
24:00 owl shift is on...

========================================================================
Follow up Notes on Previous Shift

We did the switch to BEAM_832 with ktev1 configuration from
BEAM_832_NOKTEV1.

========================================================================
Items to Follow up on Next Shift

Keep taking data with the regular BEAM_832 configuration;